Jane has $10 000 to invest. She invested part of it in a term deposit that pays 4% per year and the rest of her money was invested in bonds paying 5% per year. If Jane earned $440 in interest for one year, how much money did she invest at each rate?
Answer by checkley71(8403)

.05X+.04(10,000-X)=440
.05X+400-.04X=440
.01X=440-400
.01X=40
X=40/.01
X=4,000 AMOUNT INVESTED @ 5%.
10,000-4,000=6,000 AMOUNT INVESTED @ 4%.
PROOF:
.05*4,000+.04*6000=440
200+240=440
440=440